OpenStudy (anonymous):

can someone help me with the steps on how to solve the following proportion and solve for x?

jagr2713 (jagr2713):

Sure

OpenStudy (anonymous):

\[\frac{ 12 }{ x+3 }=\frac{ x+3 }{ x}\]

jagr2713 (jagr2713):

is there an answer choice

OpenStudy (anonymous):

no, i had to set up a proportion to find the sides of similar right triangles and i got that, but i forgot how to actually solve the proportion haha.

jagr2713 (jagr2713):

Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. x+3/x=12/x+3

jagr2713 (jagr2713):

Set up the rational expression with the same denominator over the entire equation.

jagr2713 (jagr2713):

then we multiply each term by a factor of 1 that will equate all the denominators. In this case, all terms need a denominator of x(x+3). The (x+3)/x expression needs to be multiplied by (x+3)/(x+3) to make the denominator x(x+3). The 12/(x+3) expression needs to be multiplied by (x)/(x) to make the denominator x(x+3). x+3/x⋅x+3/x+3=12/x+3⋅x/x

OpenStudy (anonymous):

wouldn't you have to cross multiply?

jagr2713 (jagr2713):

i dont do mines that way but u can idk if it works

OpenStudy (anonymous):

because thats how my teacher told us to do it. so in that case it would be \[12x=x ^{2}+9\] is that right?

jagr2713 (jagr2713):

what i got 3

OpenStudy (anonymous):

i know the answer is 3, but idk how to get that.

jagr2713 (jagr2713):

ok let me finsh

jagr2713 (jagr2713):

Multiply the expression by a factor of 1 to create the least common denominator (LCD) of x(x+3). (x+3)(x+3)x/(x+3)=12/x+3⋅x/x

OpenStudy (anonymous):

i can't see what your typing, a bunch of diamonds with ? in them are showing up

jagr2713 (jagr2713):

it changed u see it

jagr2713 (jagr2713):

Apply the distributive property. x(x+3)+3(x+3)/x(x+3)=12/x+3⋅x/x

jagr2713 (jagr2713):

Apply the distributive property. x(x)+x(3)+3(x+3)/x(x+3)=12/x+3⋅x/x

jagr2713 (jagr2713):

Apply the distributive property. x(x)+x(3)+3(x)+3(3)/x(x+3)=12/x+3⋅x/x Multiply x by x to get x2. x2+x(3)+3(x)+3(3)/x(x+3)=12/x+3⋅x/x

jagr2713 (jagr2713):

Multiply x by 3 to get x⋅3. x2+x⋅3+3(x)+3(3)/x(x+3)=12/x+3⋅x/x Move 3 to the left of the expression x⋅3. x2+3⋅x+3(x)+3(3)/x(x+3)=12/x+3⋅x/x Multiply 3 by x to get 3x. x2+3x+3(x)+3(3)/x(x+3)=12/x+3⋅x/x

OpenStudy (anonymous):

ohhh ok i got it thanks!

jagr2713 (jagr2713):

You're Welcome